OpenStudy (anonymous):

Solve for t in d = rt, if d = 57 and r = 30.

Parth (parthkohli):

\( \color{Black}{\Rightarrow 57 = 30 \times t}\) Divide both sides by 30.

OpenStudy (unklerhaukus):

\[d=rt\]divide by r\[\frac{d}{r}=\frac{rt}r\]\[\frac{d}{r}=t\]\[t=\frac {57}{30}\]